A sample of 94.9 g

94.9 g of tetraphosphorous decoxide (P4⁢O10) reacts with 68.0 g

68.0 g of water to produce phosphoric acid (H3⁢PO4) according to the following balanced equation.

P4O10+6H2O⟶4H3PO4

Determine the limiting reactant for the reaction.

Calculate the mass of H3⁢PO4 produced in the reaction.

Calculate the percent yield of H3⁢PO4 is isolated after carrying out the reaction.


1
Expert's answer
2022-03-20T06:37:30-0400

Balanced equation is given as follows


"P_4O_{10} +6H_2O \longrightarrow 4H_3PO_4"


Mass of "P_4O_{10}" = 94.9 g


Mass of H2O = 68 g


Mole of P4O10 = "\frac{94.9}{283.88}= 0.334"


Mole of "H_2O= \frac{68}{18}=3.78"


1 mole of P4O10 reacts with 6 mole of H2O to give 4 mole of H3PO4


so, 0.334 mole of P4O10 reacts with 2.004 mole of H2O


so here limiting reagent is P4O10


mole of H3PO4 produces = 1.336 mole


mass of H3PO4 produces = 130.9 g


% yield of the reaction is "\frac{1.336 }{4}\times 100=33.4 %" %